OpenStudy (anonymous):

Determine which system below will produce infinitely many solutions. Choices: -4x + 6y = 20 and 4x - 3y = -2 -2x + 4y = 18 and 4x - 8y = -36 3x - 2y = 5 and -6x - 4y = -10 5x - 3y = 12 and 15x - 9y = 24

OpenStudy (amistre64):

which one is just a multiple of the other?

OpenStudy (anonymous):

what do you mean?

OpenStudy (amistre64):

i mean, given ax + by = c, which one shows the other equation for some multiple n: an x + bn y = cn

OpenStudy (amistre64):

for example: 5x - 3y = 12 15x - 9y = 24 is one a multiple of the other? is there some value n, that we can mutliply by that will turn one into the other? 5x - 3y = 12 *3 *3 *3 15x - 9y = 24 ^^ this doesnt work 12*3 = 36

OpenStudy (anonymous):

Oh okay so you just do that to find out which is one multiple of the other and which ever one is that's the answer?

OpenStudy (amistre64):

another trick is to divide one by the other and see if they all form the same value lets try #1, stack them one over the other and divide the numbers -4x + 6y = 20 4x - 3y = -2 -------------- -1 -2 -10 these arent all the same so its not this one either

OpenStudy (amistre64):

yes, which ever one fits the bill isthe solutions

OpenStudy (amistre64):

dividing using the first example: 15x - 9y = 24 5x - 3y = 12 ------------- 3 3 2 notice we still get a bad result and therefore its not this one still

OpenStudy (anonymous):

Alright I think I got what you meant on how to do this :) thank you.

OpenStudy (amistre64):

youre welcome :)